OpenStudy (anonymous):

Does {(1,3),(3,4),(2,3)}span (Z/5Z)^2? Z=integer

OpenStudy (anonymous):

I don't understand (again) your notation: (Z/5Z)^2 means what?

jimthompson5910 (jim_thompson5910):

Z/5Z is the integers mod 5 right?

OpenStudy (anonymous):

yeaps

jimthompson5910 (jim_thompson5910):

best way to do this is to make a table of all the possible ordered pairs mod 5 and see if you can make each value in the table using only the given ordered pairs

jimthompson5910 (jim_thompson5910):

if you can, then that set spans (Z/5Z)^2

OpenStudy (anonymous):

but what does the power 2 mean though? is there any significance?

jimthompson5910 (jim_thompson5910):

Z^2 is the set of all ordered pairs drawn from Z

jimthompson5910 (jim_thompson5910):

(Z/5Z)^2 is the set of all ordered pairs drawn from Z mod 5 = {0, 1, 2, 3, 4}

jimthompson5910 (jim_thompson5910):

more formally, \[\large \frac{\mathbb{Z}^2}{5\mathbb{Z}}=\{(a,b)|a,b\in\frac{\mathbb{Z}}{5\mathbb{Z}}\}\]

OpenStudy (anonymous):

hmm okay. i think i get it now. thanks!

jimthompson5910 (jim_thompson5910):

oh and you might see it as \[\large \mathbb{Z}^2_{5}=\{(a,b)|a,b\in\mathbb{Z}_{5}\}\]

jimthompson5910 (jim_thompson5910):

If you're still stuck, notice that (1,3)+(3,4)=(4,7)=(4,2) and (4,2)+(1,3)=(5,5)=(0,0) <--- this is one key ingredient Also, (4,2)+(2,3)=(6,5)=(1,0) <--- another useful (very useful) element Finally, (3,4)+(2,3)=(5,7)=(0,2) and (0,2)+(0,2)+(0,2)=(0,6)=(0,1) ----------------------------------------- Why did I do all that? Well for one thing, I've shown that (0,0) is in there. This is the additive identity. Also, I've shown that (1,0) and (0,1) can also be generated So we can then say that any element x can be generated by the linear combination x = c*(1,0)+d*(0,1) to get EVERY element thats in the desired set that we want. So therefore, {(1,3),(3,4),(2,3)} spans (Z/5Z)^2

jimthompson5910 (jim_thompson5910):

ok now that I think about it, you don't need to generate (0,0) explicitly since the formula involving x takes care of it.